# A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v2

@@ Abstract
Poisoning efficiency plays a critical role in poisoning-based backdoor attacks. To evade detection, attackers aim to use the fewest poisoning samples while achieving the desired attack strength. Although efficient triggers have significantly improved poisoning efficiency, there is still room for further enhancement. Recently, selecting efficient samples has shown promise, but it often requires a proxy backdoor injection task to identify an efficient poisoning sample set. However, the proxy attack-based approach can lead to performance degradation if the proxy attack settings differ from those used by the actual victims due to the shortcut of backdoor learning. This paper presents a Proxy attack-Free Strategy (PFS) designed to identify efficient poisoning samples based on individual similarity and ensemble diversity, effectively addressing the mentioned concern. The proposed PFS is motivated by the observation that selecting the to-be-poisoned samples with high similarity between clean samples and their corresponding poisoning samples results in significantly higher attack success rates compared to using samples with low similarity. Furthermore, theoretical analyses for this phenomenon are provided based on the theory of active learning and neural tangent kernel. We comprehensively evaluate the proposed strategy across various datasets, triggers, poisoning rates, architectures, and training hyperparameters. Our experimental results demonstrate that PFS enhances backdoor attack efficiency, while also exhibiting a remarkable speed advantage over prior proxy-dependent selection methodologies.
@@
11footnotetext: Equal Contribution.
@@ I Introduction
The abundance of training data plays a vital role in the success of Deep Neural Networks (DNNs). For instance, GPT-3 [3], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pretraining on a massive corpus of 45 TB of text data. Responding to the increasing demand for data, both users and businesses are increasingly turning to third-party sources or online repositories, which offer more convenient way for data collection. However, recent researches [4, 14, 13] have demonstrated that such practices can be maliciously exploited by attackers to poison training data, thereby negatively impacting the functionality and reliability of trained models.
@@
During the training phase, a significant threat emerges in the form of backdoor attacks [4, 14, 33]. This attack involves injecting a covert backdoor into the DNNs by introducing a small number of poisoning samples into an otherwise benign training dataset. While the model appears normal when presented with benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the objective of attackers. As research continues to reveal this threat in various tasks such as speaker verification [59] and malware detection [24], attention to backdoor attacks is increasing.
Poisoning efficiency stands as a pivotal metric for attackers engaging in backdoor attacks against DNNs [60, 26, 50, 52]. Their objective revolves around poisoning the training dataset with minimal poisoning samples while achieving the desired outcomes. An ideal scenario is one where a single poisoning sample suffices to implant a backdoor, greatly enhancing the stealthiness of the attack. Currently, strategies aimed at enhancing poisoning efficiency fall into two main categories: designing efficient triggers [60, 61, 57] and selecting efficient samples [50]. The former focuses on identifying trigger patterns that are easier for models to learn, while the latter centers around the careful selection of representative samples for poisoning. This paper aligns with the latter approach, exploring the differences in poisoning efficiency of different poisoning samples in backdoor attacks.
@@
A recent investigation [50] into sample selection has yielded significant results, demonstrating that an equivalent attack success rate can be achieved using only $47\%$ to $75\%$ of the poisoning samples in comparison to random selection strategies. However, it has become apparent that this method relies on a proxy backdoor task to assess the contribution of each poisoning sample, thus constructing an efficient poison set closely associated with the specific proxy attack of the attacker. This proxy attack-based selection method [50] poses a challenge due to potential deterioration in its effectiveness when the proxy attack settings of attackers deviate from the actual settings of the victims—a frequent attack scenario [60, 61, 57]. Moreover, conducting the proxy attack task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
@@
Figure 1: The pipeline of poisoning-based backdoor attacks typically involves an attacker who combines a clean dataset (D) with a trigger (T) to create a poisoned dataset (D+T), which is then released to the victims. The victims download the poisoned data and use it to train their DNN models, applying various data transformations and augmentations (F) and hyperparameters (H)III during training. As a result, the DNN models can be infected with the backdoor trigger, which can be activated by a specific trigger condition during the inference phase.
This paper presents a solution to the aforementioned challenge by introducing a novel Proxy attack-Free Strategy (PFS) that not only achieves higher backdoor attack efficiency but also boasts a remarkable speedup compared to previous selection methods. Our contributions can be summarized as follows:
@@
  • This paper makes a comprehensive study of the forensic properties inherent in efficient poisoning samples within the context of backdoor attacks. In particular, we delve into two critical facets: the similarity between clean samples and their corresponding poisoning samples, and the diversity of poisoning set.
@@
  • Drawing on the insights derived from our analysis, we introduce a novel selection approach named the Proxy attack-Free Strategy (PFS). Distinguishing itself from previous study, PFS eliminates the requirement for a proxy attack task, effectively addressing the challenge arising from differing settings while significantly reducing the time overhead.
@@
  • The effectiveness of PFS is subjected to evaluation across three diverse datasets: CIFAR-10, CIFAR-100, and Tiny-ImageNet. Our empirical findings consistently confirm its efficacy. It is worth noting that PFS has achieved superior attack success rates while achieving a remarkable acceleration when contrasted with prior proxy-dependent selection methodologies. This attribute renders PFS particularly well-suited for scalability, making it apt for accommodating larger datasets and complex models.
@@ II-A Backdoor Attacks
Designing Efficient Triggers. Trigger design is a hot topic in backdoor learning. Chen et al. [4] initially proposed a blended strategy for evading human detection, acquiring poisoning samples by blending clean samples with triggers. Subsequent studies have explored leveraging natural patterns such as warping [38], rotation [48], style transfer [5], frequency [9, 58], and reflection [34] to create triggers that are more imperceptible and efficient compared to previous methods. Drawing inspiration from Universal Adversarial Perturbations (UAPs) [37] in adversarial examples, some research [61, 25, 7] optimizes an UAP on a pre-trained clean model as the trigger, a highly effective and widely adopted approach. Unlike previous approaches that use universal triggers, Li et al. [27] employ GANs models to generate sample-specific triggers.
Selecting Efficient Poisoning Samples. Efficient poisoning samples selection is an under-explored area and is orthogonal to the trigger design. Xia et al. [50] first explored the contribution to backdoor injection for different data. Their work illustrated the inequitable influence of individual poisoning samples and underscored the enormous potential for improved data efficiency within backdoor attacks through efficient sample selection. Simultaneously, Gao et al. [10] assert a similar notion—that not all samples equally foster the process of poisoning in backdoor attacks. They introduce the concept of ”forgetting events” to spotlight the most potent poisoning samples in the context of clean-label backdoor attacks. However, both of these studies [50, 10] rely on proxy attacks to determine the efficient subset of samples. Regrettably, this dependence on proxy attacks introduces a vulnerability, leading to a degradation in attack performance when disparities arise in parameters like Transformation (F) and Hyperparameters (H) between the proxy poisoning attack and the actual poisoning process, as depicted in Figure 1. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper. Comparable to our approach, the Outlier Poisoning Strategy (OPS) [17] also utilizes a surrogate model for efficient sample selection in backdoor attacks against anti-spoof rebroadcast detection. The primary concept behind this strategy involves embedding triggers into samples that pose the greatest classification challenge (referred to as outlier samples), encouraging the network to depend on the trigger for classification. However, OPS [17] is specifically designed for binary classification and clean-label settings, exhibiting diminished performance in scenarios involving dirty-label settings and multiple classifications.
@@ II-B Backdoor Defenses
Backdoor attacks pose a significant threat to the security of Deep Neural Network (DNN), leading to the emergence of various defense mechanisms aimed at mitigating these risks. The current landscape of backdoor defenses can be broadly categorized into two classes: backdoor detection and backdoor erasure. Backdoor detection aims to identify the presence of poisoning either in input data or within the employed DNN models themselves [11, 18, 8, 53]. Conversely, backdoor erasure strategies aim to repair poisoning models by removing the embedded backdoor, either during or after the training process. In the training phase, the concept of training clean models on poisoned data has been introduced, exemplified by Li et al. [28]. Post-training, a range of techniques, including fine-tuning, distillation, and pruning, can be employed to eliminate the backdoor from the poisoning model [29, 47].
III Degradation of proxy attack-based Selection
The purpose of this section is to emphasize the limitations associated with sample selection methods that rely on proxy attacks, which serves as a motivation for designing a proxy attack-free sample selection strategy in backdoor attacks. To better understand our key observations, we will first review the pipeline of backdoor attacks and sample selection methods based on proxy attack.
Poisoning-based Backdoor Attacks. Poisoning-based backdoor attacks involve injecting a backdoor into a subset $\mathcal{P^{\prime}}$ of a clean training set $\mathcal{D}=\{(x_{i},y_{i})|i=1,\cdots,N\}$, resulting in a corresponding poisoned set $\mathcal{P}=\{(x^{\prime}_{i},k)|x^{\prime}_{i}=\mathcal{T}(x_{i},t),(x_{i},y_{i})\in\mathcal{P^{{}^{\prime}}},i=1,\cdots,P\}$. Here, $\mathcal{T}(x,t)$ is a pre-defined generator that adds the trigger $t$ into clean sample $x$, while $y_{i}$ and $k$ represent the true label and attack-target label of the clean sample $x_{i}$ and the poisoning sample $x^{\prime}_{i}$, respectively. After poisoning, the victim will train a deep model:
@@ III Degradation of proxy attack-based Selection
In the above equation, $f_{\theta}$ represents the DNN model, and $L$ is the corresponding loss function. The poisoning rate $r$ can be calculated as $r=\frac{P}{N}$. The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set $\mathcal{P}$. This approach is orthogonal to the mainstream method of improving poisoning efficiency and has not been thoroughly explored yet.
Proxy Attacks-based Sample Selection. Recent investigations [50, 55, 10] have highlighted that diverse samples contribute differently to the backdoor injection and a particular subset with efficient data has higher attack success rate. Notably, [50, 10] claim that the efficacy of poisoning samples is intertwined with instances of ”forgetting events” during the backdoor injection process. To distill the most efficient subset, [50]***It is worth noting that both [50] and [10] employ the concept of forgetting events to guide the selection of poisoned sets. It is worth highlighting that while FUS [50] employs multiple iterations of screening, [10] adopts a one-step screening approach, rendering [10] a specific case of FUS. As such, for the purpose of analyzing the drawbacks of proxy attack-based methods for selecting poisoning samples, we treat FUS as a representative example in this section. formulates the sample selection as an optimization problem and introduce a Filtering-and-Updating Strategy (FUS). Although FUS has acquired remarkable performance, it uses a proxy attack processing to find the efficient subset.
As illustrated in Figure 1, poisoning-based backdoor attacks contain two phases: The attacker combines data (D) and trigger (T) to build poisoned data (D+T) and release it. Victims download the released data and adopt it to train infected DNN models. Notably, during the training process, various data transformations/augmentations (F) and hyperparameters (H)In this context, H encompasses various training settings, encompassing elements like the model architecture, optimizer choice, and training hyperparameters. have been applied by victims. Generally, the attacker has no access to any information of the victim phase and only performs the poisoning injection at the attack phase. However, proxy attack-based sample selection methods need to select samples with the help of the complete attack process (both attacker and victim phases). We argue that all parts involved in backdoor attacks are the factors that affect the efficiency of poisoning samples, containing D, T, F, and H. When there contains some gaps (e.g. Transform (F) and Hyperparameters (H) in Figure 1) between proxy poisoning attack and actual poisoning process of victims, the poisoning effectiveness of the selected samples may be decreased. Following, we will explore what will happen if the actual poisoning process is different from the proxy poisoning attack in proxy attack-based sample selection methods.
III-A Experimental Settings
To investigate the proposed question within this section, we engage in a series of experiments utilizing the VGG16 model [43] as the victim model. These experiments are conducted on the CIFAR-10 dataset [23]. Our chosen config encompasses the SGD optimizer, initially set with a learning rate of 0.01, a momentum value of 0.9, and a weight decay coefficient of 5e-4. Notably, we incorporate two widely adopted data transformations—random cropping and random horizontal flipping. The training is performed for 70 epochs with a batch size of 256. Poisoning samples $x^{\prime}_{i}$ are generated using a blending strategy [4] in which the clean samples $x_{i}$ are blended with a trigger image $t$ using the equation $x^{\prime}_{i}=\mathcal{T}(x_{i},t)={{\lambda}{\cdot}{t}}+(1-{\lambda}){\cdot}{x_{i}}$, where ${\lambda}$ is set to 0.15. We set category 0 as the target for the attack, and we set the poisoning rate with $r=0.01$—equating to a poisoning set size of $P=500$. The process involves 15 iterations of selection, aligning with the settings of [50]. In accordance with previous literature [26], the attack success rate is defined as the probability of classifying a poisoned test data to the target label. Unless stated otherwise, all experiments in this section follow the aforementioned settings. To explore the degeneration in poisoning efficiency within selected samples due to disparities between the proxy poisoning attack and the actual poisoning process of victims, we compare the Filtering-and-Updating Strategy (FUS) with SELF-FUS. SELF-FUS represents an ideal scenario for FUS, assuming the attacker possesses complete information about both the attacker and victim phases during poisoning-based backdoor attacks—a scenario that doesn’t align with the threat model in this study. Therefore, proxy poisoning attack of attacker phase is consistent with the actual poisoning process of victim phase in SELF-FUS, which is not guaranteed in the FUS.
III-B Empirical Studies
Impact of Different Data Transformations (F) on Proxy Attack-based Sample Selection Efficacy. We investigated how different data transforms (F) affect the effectiveness of proxy attack-based sample selection methods. To conduct our analysis, we used the experimental settings described previously and followed them exactly. Specifically, we configured the data transformations for the proxy poisoning attack during the sample selection phase (FUS) as ”RandomCrop” and ”RandomHorizontalFlip”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of transformations, including ”None”, ”RandomCrop”, ”RandomHorizontalFlip”, ”RandomRotation”, and ”ColorJitter”. Our results, depicted in Figure 2, indicate that different data transforms (F) decrease the effectiveness of proxy attack-based sample selection methods. An exception to this pattern emerges in cases where the settings for both the proxy and actual attacks align closely, exemplified by the ”1+2” searching and ”1” attacking scenario. This particular outcome is logical and can be attributed to the inherent margin of error.
Figure 3: The ASR in scenarios where the hyperparameters (H) used during the proxy poisoning attack within attacker phase in FUS is different from the actual poisoning process within victim phase. The coordinates ”V-13, SGD, 0.01”III-B on the horizontal axis symbolize the utilization of VGG-13 as the architecture and SGD with an initial learning rate of 0.1 as the optimizer. Each reported result is an average computed from five separate runs.
Figure 4: The attack success rate in situations where both the data transformation (F) and the hyperparameter (H) used during the actual attack within the victim phase is different from that used within the search process with FUS [50] on CIFAR-10 and Tiny-ImageNet datsets. The horizontal coordinates are labeled as ”0”, ”1”, ”2”, ”3”, and ”4”, representing different combinations of data transforms and hyperparameters. All results are computed as the mean of five different runs.
Effect of Different Hyperparameters (H) on Proxy Attack-based Sample Selection Efficiency. In this section, We thoroughly investigated the impact of different hyperparameters on the effective data efficiency searched by FUS [50]. Adhering to the aforementioned experimental framework, we configured the hyperparameter combination for the proxy poisoning attack during the sample selection phase as ”V-16, SGD, 0.01”. Subsequently, during the actual poisoning attack within the victim phase, we introduced variations by employing a range of hyperparameters, including ”V-13, SGD, 0.01”, ”V-13, Adam, 0.001”, ”V-16, Adam, 0.001”, ”R-18, SGD, 0.01”, ”R-18, Adam, 0.001”, and ”P-18, SGD, 0.01”In this context, V-13, V-16, R-18, and P-18 denote VGG13, VGG16, ResNet18, and PreActResNet18, respectively. . Figure 3 indicates that different hyperparameters (H) decrease the effectiveness of proxy attack-based sample selection methods. Once again, we observed anomalous occurrences when the parameters for both proxy searching and actual attacks closely converged, as illustrated by the scenario involving ”V-16-SGD-0.01” proxy searching and subsequent ”V-13-SGD-0.01” attacks within the victim phase.
Summary Collectively, our findings underscore the susceptibility of proxy attack-based sample selection techniques to shifts in data transforms and hyperparameters deployed during the victim phase. Notably, SELF-FUS consistently outperforms FUS across various scenarios, with the exception of outliers observed in both Figure 2 and Figure 3. These outliers, marked by the ”1+2” searching and ”1” attacking in Figure 2, as well as the ”V-16-SGD-0.01” searching and ”V-13-SGD-0.01” attacking in Figure 3, share closely aligned proxy and actual attack settings. To further validate our hypothesis, we continue to simulate more realistic attack scenarios that introduce disparities in both data transformations (F) and hyperparameters (H) between proxy and actual attacks. Figure 4 shows the ASR comparison between FUS and SELF-FUS on both CIFAR-10 and Tiny-ImageNet datasets. The notable difference in ASR between FUS and SELF-FUS underscores the significant impact of proxy attack-based methods on sample efficiency. Specifically, it highlights how FUS experiences serious degradation in sample efficiency when there is a substantial difference between the settings for proxy poisoning attacks during the attacker phase and those employed in the actual poisoning process during the victim phase.
IV Methodology
To improve poisoning efficiency, it is important to develop proxy attack-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we will first discuss some key observations.
@@ IV Methodology
IV-A Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks

### v3

@@ Abstract
Poisoning efficiency is crucial in poisoning-based backdoor attacks, as attackers aim to minimize the number of poisoning samples while maximizing attack efficacy. Recent studies have sought to enhance poisoning efficiency by selecting effective samples. However, these studies typically rely on a proxy backdoor injection task to identify an efficient set of poisoning samples. This proxy attack-based approach can lead to performance degradation if the proxy attack settings differ from those of the actual victims, due to the shortcut nature of backdoor learning. Furthermore, proxy attack-based methods are extremely time-consuming, as they require numerous complete backdoor injection processes for sample selection. To address these concerns, we present a Proxy attack-Free Strategy (PFS) designed to identify efficient poisoning samples based on the similarity between clean samples and their corresponding poisoning samples, as well as the diversity of the poisoning set. The proposed PFS is motivated by the observation that selecting samples with high similarity between clean and corresponding poisoning samples results in significantly higher attack success rates compared to using samples with low similarity. Additionally, we provide theoretical foundations to explain the proposed PFS. We comprehensively evaluate the proposed strategy across various datasets, triggers, poisoning rates, architectures, and training hyperparameters. Our experimental results demonstrate that PFS enhances backdoor attack efficiency while also offering a remarkable speed advantage over previous proxy attack-based selection methodologies.
@@
22footnotetext: Corresponding Author.
@@ I Introduction
The abundance of training data is crucial to the success of Deep Neural Networks (DNNs) [39]. For example, GPT-3 [5], a deep learning model with 175 billion parameters, owes its effectiveness in various language processing tasks to its pre-training on a massive corpus of 45 TB of text data. In response to the increasing demand for data, users and businesses are increasingly turning to third-party sources or online repositories for more convenient data collection. However, recent researches [6, 17, 16, 3, 8, 4, 41] have demonstrated that these practices can be maliciously exploited by attackers to poison training data, thereby negatively impacting the functionality and reliability of trained models.
@@
During the training phase, a significant threat emerges in the form of backdoor attacks [6, 17, 45, 38]. These attacks involve injecting a covert backdoor into DNNs by introducing a small number of poisoned samples into an a benign training dataset. While the model appears normal when presented with benign inputs, a predefined trigger can activate the infected model, forcing its predictions to align with the attackers’ objectives. As research continues to reveal this threat in various tasks, such as speaker verification [74], malware detection [29], emotion analysis [71, 42], and Deepfake detection [56], attention to backdoor attacks is increasing.
@@
Recent studies have sought to enhance the efficiency of backdoor attacks by selecting optimal poisoning samples. These studies [63, 13, 21, 40] demonstrate that an equivalent attack success rate can be achieved using only 47% to 75% of the poisoning samples compared to random selection strategies. However, it has become apparent that these methods rely on a proxy backdoor task to assess the contribution of each poisoning sample, thus constructing an efficient poison set closely associated with the specific proxy attack of the attacker. These proxy attack-based selection methods [63, 13, 21] pose a significant challenge due to potential deterioration in its effectiveness when the proxy attack settings of attackers deviate from the actual settings of the victims—a frequent scenario in attacks [75, 76, 72]. Moreover, conducting the proxy attack task significantly increases the execution time of the algorithm, limiting its scalability on large datasets.
@@
Figure 1: The pipeline of poisoning-based backdoor attacks typically involves an attacker who combines a clean dataset (D) with a trigger (T) to create a poisoned dataset (D+T), which is then released to the victims. The victims download the poisoned data and use it to train their DNN models, applying various data transformations and augmentations (F) and hyperparameters (H)*IV during training. As a result, the DNN models can be infected with the backdoor trigger, which can be activated by a specific trigger condition during the inference phase.
This paper addresses the aforementioned challenges by introducing a novel Proxy attack-Free Strategy (PFS). Unlike previous methods reliant on proxy attack tasks, PFS leverages similarity measures between benign and corresponding poisoning samples. Specifically, we utilize a pre-trained feature extractor to compute cosine similarity within the feature space. We demonstrate that selecting samples with high similarity between benign and corresponding poisoning samples significantly enhances attack success rates compared to low-similarity samples. However, solely using high-similarity samples for poisoning resulted in lower success rates compared to random sampling. We attribute this performance decline to the overly constrained diversity within the poisoning sample pool. Therefore, our paper emphasizes both similarity between benign and corresponding poisoning samples and diversity within poisoning samples as critical factors influencing sample efficiency in backdoor attacks. Additionally, we provide theoretical foundations to elucidate this phenomenon. Based on these insights, we propose PFS, which incorporates a pre-trained feature extractor to filter out inefficient samples based on cosine similarity and integrates random selection to enhance sample diversity within the poisoning set. We evaluate this strategy across CIFAR-10, CIFAR-100, and Tiny-ImageNet datasets. Our experimental results demonstrate that PFS significantly improves backdoor attack efficiency while offering a notable speed advantage over proxy attack-based methodologies. Contributions can be summarized as follows:
@@
  • We conduct a comprehensive study of the forensic properties inherent in efficient poisoning samples within the context of backdoor attacks.
@@
  • Based on the insights from our analysis, we introduce a novel selection approach named the Proxy attack-Free Strategy (PFS). Unlike previous studies, PFS eliminates the requirement for a proxy attack task, effectively addressing challenges arising from differing settings and significantly reducing the time overhead.
@@
  • The proposed PFS has achieved superior attack success rates while achieving a remarkable acceleration when contrasted with prior proxy attack-based selection methodologies. This attribute renders PFS particularly well-suited for scalability, making it apt for accommodating larger datasets and complex models.
@@ II-A Backdoor Attacks
Designing Efficient Triggers. Trigger design is a hot topic in backdoor learning. Chen et al. [6] initially proposed a blended strategy for evading human detection, acquiring poisoning samples by blending clean samples with triggers. Subsequent studies have explored leveraging natural patterns such as warping [50], rotation [61, 67], style transfer [7], frequency [12, 73], and reflection [46] to create triggers that are more imperceptible and efficient compared to previous methods. Drawing inspiration from Universal Adversarial Perturbations (UAPs) [49] in adversarial examples, some research [76, 30, 10] optimizes an UAP on a pre-trained clean model as the trigger, a highly effective and widely adopted approach. Unlike previous approaches that use universal triggers, Li et al. [33] employ GANs models to generate sample-specific triggers.
Selecting Efficient Poisoning Samples. Efficient poisoning samples selection is an under-explored area and is orthogonal to the trigger design. Xia et al. [63] first explored the contribution to backdoor injection for different data. Their work illustrated the inequitable influence of individual poisoning samples and underscored the enormous potential for improved data efficiency within backdoor attacks through efficient sample selection. Simultaneously, Gao et al. [13] assert a similar notion—that not all samples equally foster the process of poisoning in backdoor attacks. They introduce the concept of ”forgetting events” to spotlight the most potent poisoning samples in the context of clean-label backdoor attacks. Additionally, the RD score [62] measures the distance between the model’s output and the target class, enabling the selection of the most impactful samples for poisoning. To further refine this process, the authors employ a greedy search algorithm that iteratively selects samples with the highest RD scores. However, all of these studies [63, 13, 62] rely on proxy attacks to determine the efficient subset of samples. Regrettably, this dependence on proxy attacks introduces a vulnerability, leading to a degradation in attack performance when disparities arise in parameters like Transformation (F) and Hyperparameters (H) between the proxy poisoning attack and the actual poisoning process, as depicted in Figure 1. Therefore, a sample selection strategy that is not based on proxy attacks is urgently needed, and is the main contribution of our paper. Comparable to our approach, the Outlier Poisoning Strategy (OPS) [21] also utilizes a surrogate model for efficient sample selection in backdoor attacks against anti-spoof rebroadcast detection. The primary concept behind this strategy involves embedding triggers into samples that pose the greatest classification challenge (referred to as outlier samples), encouraging the network to depend on the trigger for classification. However, OPS [21] is specifically designed for binary classification and clean-label settings, exhibiting diminished performance in scenarios involving dirty-label settings and multiple classifications.
Recently, a few works also exploited backdoor for positive purposes [31, 37, 20, 68], which are out of our scope.
@@ II-B Backdoor Defenses
Backdoor attacks pose a significant threat to the security of Deep Neural Networks (DNNs), leading to the emergence of various defense mechanisms aimed at mitigating these risks. The current landscape of backdoor defenses can be broadly categorized into two classes: backdoor detection and backdoor erasure. Backdoor detection aims to identify the presence of poisoning either in input data or within the employed DNNs themselves [14, 22, 11, 65, 26, 28]. Conversely, backdoor erasure strategies aim to repair poisoning models by removing the embedded backdoor, either during or after the training process. In the training phase, the concept of training clean models on poisoned data has been introduced by [34]. Post-training, a range of techniques, including fine-tuning, distillation, and pruning, can be employed to eliminate the backdoor from the poisoning model [35, 60].
III Methodology
To improve poisoning efficiency, it is important to develop proxy attack-free sample selection methods that leverage the forensic features of efficient data in poisoning-based backdoor attacks. Before introducing our method, we first review the pipeline of backdoor attacks and discuss some observations.
III-A Poisoning-based Backdoor Attacks.
Poisoning-based backdoor attacks involve injecting a backdoor into a subset $\mathcal{P^{\prime}}$ of a clean training set $\mathcal{D}=\{(x_{i},y_{i})|i=1,\cdots,N\}$, resulting in a corresponding poisoned set $\mathcal{P}=\{(x^{\prime}_{i},k)|x^{\prime}_{i}=\mathcal{T}(x_{i},t),(x_{i},y_{i})\in\mathcal{P^{{}^{\prime}}},i=1,\cdots,P\}$. Here, $\mathcal{T}(x,t)$ is a pre-defined generator that adds the trigger $t$ into clean sample $x$, $y_{i}$ and $k$ represent the true and attack-target labels of the clean sample $x_{i}$ and the poisoning sample $x^{\prime}_{i}$, respectively. After poisoning, victim will train a deep model:
@@ III Methodology
In the above equation, $f_{\theta}$ represents the DNN model, and $L$ is the corresponding loss function. The poisoning rate $r$ can be calculated as $r=\frac{P}{N}$. The goal of sample selection methods is to enhance the efficiency of backdoor attacks by selecting an optimal poisoning set $\mathcal{P}$.
@@
III-B Forensic Features of Efficient Data in Poisoning-based Backdoor Attacks